OpenStudy (anonymous):

Choose the equation of the line passing through the point (2, 6) and parallel to y = -3x - 4. y = 3x - 12 y = -3x y = -3x + 12 y = 3x

OpenStudy (anonymous):

easy, y= -3x + B Than just plaggin the given point:)

OpenStudy (anonymous):

y= -3x +B 6 = -3(2)+B 6= -6 +B B= 12 y=-3x +12

OpenStudy (anonymous):

again, use the (y-y1)=m(x-x1)

OpenStudy (anonymous):

Yeah, you can also use that formula. However I like the other format is very simple. If the slopes are PARALLEL than you keep the slope the same from the original equation and plug in whatever order pair/solution they give you. If they want Perpendicular than you take the original slope and you take the negative reciprical of it. Ex. y= -3X Parallel: y= -3x + B; plug in (x,y) to ge B and rewrite final equation. Perpendicular: y= 1/3 x + B

OpenStudy (anonymous):

thanks guise!